# Modeling lower-truncated and right-censored insurance claims with an extension of the MBBEFD class

### v1

@@ 3.2 Properties of the MBBEFD class
This derivative is symmetric around zero. We have the following result.
@@
scaled by a constant factor $(1+a)\log(1/b)>0$. It is symmetric around the mode $z^{*}$, and it decays slower than the Gaussian density. This now shows why the MBBEFD densities are not sufficient for general insurance claims modeling, because general insurance claims are typically positively skewed, which cannot be captured by the logistic density.
@@ 3.3 Extension of Bernegger’s idea using non-bell shaped densities
We extend the class of bell-shaped MBBEFD densities to a more general class of exposure curves, which allows, in particular, for skewness in their corresponding densities. We call this extended family the Bernegger class. In Section 2.1, we have started from a generic exposure curve $G:[0,1]\to[0,1]$ which is a non-decreasing, concave and twice continuously differentiable function with the normalization $G(0)=0,\,G(1)=1$ and with $G^{\prime}(0)>0$. The MBBEFD exposure curve (3.1) can be reparametrized. Indeed, by using (3.8), we obtain for $z\in[0,1]$
@@
for $b\geq 0$ and $a>-\min(1,b)$; we refer to Section 3.1 of Bernegger [2]. This structure can be used to design exposure curve forms that do not have the bell-shape property of Proposition 3.3. The function $a+b^{z}=a+\exp\{-z\log(1/b)\}$ is strictly decreasing and convex for $b\in(0,1)$. We modify the modeling set-up (3.9) as follows. Choose a function $B:[0,1]\rightarrow{\mathbb{R}}$ that satisfies
@@
which ensures that the normalization property $G(0)=0$ and $G(1)=1$ is satisfied. The function $h$ will be denoted as the link function, whereas the function $b$ will be named the inner function, and we notice that Bernegger’s original choice was $b(z)=a+b^{z}$ for $b\in(0,1)$ and $h(x)=\log(x)$, meaning that he used a logarithmic linked exposure curve. We will first explore some examples using the same link function and then introduce the exponentially linked exposure curves, which use the link function $h(x)=\exp(x)$ and seem to be more suitable in modeling general insurance claims. We call the class of distributions induced by exposure curves of the form (3.11) the Bernegger class.
@@ 4 Logarithmic linked exposure family
The function $G$ is an exposure curve if and only if one of the following holds:
@@
For general insurance pricing, we are interested into unimodal densities and one is thus interested in the first derivative of $f_{Z}$ in order to characterize the maximum of the density
@@
Note that this derivative only exists if the third derivative of $b$ exists.
@@ 5 Exponentially linked exposure family
The function $G$ is an exposure curve if and only if one of the following holds:
@@
As for the logarithmic linked exposure curves, one can then derive using Theorem 2.1 a distribution function leading to an absolutely continuous density on $[0,1)$ and to a point mass in 1.
@@
Assume that a twice continuously differentiable function $b:[0,1]\to(0,\infty)$ with $b(0)\neq b(1)$ fulfills condition (5.2) or (5.3). The exposure curve $G$ defined in (5.1) provides the distribution of a random variable $Z\sim F_{Z}$
@@
The first derivative of the density $f_{Z}$ given in (5.5) allows to characterize its extrema and is given by
@@
Note that this derivative only exists if the third derivative of $b$ exists.
@@
A well-known example of a distribution belonging to the exponentially linked exposure family arises when choosing $b(z)=-\lambda z$ for $z\in[0,1]$ and $\lambda>0$. The distribution obtained using Corollary 5.2 is the distribution of $Z=\min\{Y,1\}$, where $Y\sim\textrm{Exp}(\lambda)$. Indeed, the construction in (5.4) gives as distribution function
$F_{Z}(z)=\left(1-e^{-\lambda z}\right)\mathds{1}_{\{z<1\}}+\mathds{1}_{\{z=1\}},$
for $z\in[0,1]$. $\blacksquare$
6 Examples
In this section, the goal is to explore some examples belonging to the logarithmic and the exponentially linked exposure family of the Bernegger class. These examples will be used to fit general insurance claims data using the tractability of the models described in Section 2.1. In other words, MLE can directly be used since the distribution functions as well as their associated densities are of closed form.
For this, we will vary the choice of the inner function $b(z)$. In the following, $\theta\in\Theta$ will denote the set of parameters appearing in the inner function. Using a dataset for $Z$ taking values in $[0,1]$, we will fit the models with MLE, maximizing the log-likelihood function
@@ 6 Examples
Figure 3: Histogram (lhs) and empirical density (rhs) of the claims $Z_{i}$.
These normalized losses $Z_{i}$, along with different models, will be used to solve the above MLE maximization problems in order to produce the results of this section. For this, we use the R function optim in order to minimize the sum of the negative of the log-likelihoods evaluated at $Z_{i}$, after having possibly transformed our parameters in a way to ensure that they lie in a suitable open domain, as described below. The first fitted model is the classical MBBEFD model of Bernegger [2].
@@ 6.1 The MBBEFD example
We have seen in Section 3.3 that the MBBEFD example belongs to the logarithmic linked exposure family and is obtained by choosing
@@
for parameters $b\geq 0$ and $a>-\min(1,b)$. Using the parametrization in (3.8), it is possible to give the conditions under which this class of distributions leads to unimodal densities on $[0,1)$. Indeed, according to Proposition 3.2, such unimodal densities are obtained if and only if
@@
in order to find unimodal MLE solutions of the standard and extended maximization problems described in (6.1) and (6.2). In Figure 4, the empirical density of the data points strictly smaller than 1 is plotted in blue color. Since this corresponds to a true density integrating to one, we further show the conditional density of $Z|_{\{Z<1\}}$ of the fitted models, using the parameters obtained by solving the standard problem (6.1) (in green) and the extended problem (6.2) (in red). The point mass and the mean are shown in Table 1, as well as the log-likelihoods of the random variables $Z|_{\{Z<1\}}$ and $Z$, and the AIC scores that are computed using $\ell_{Z}$. We see that, as expected, the MLE solution of the extended problem gives better results, even if the densities obtained are not close to the empirical density. Proposition 3.3 helps us to understand why the fit is not accurate since the empirical density is skewed to the right, whereas the MBBEFD class only allows for symmetric densities.
Figure 4: The MBBEFD: densities of the random variable $Z|_{\{Z<1\}}$.
@@
Table 1: The MBBEFD: results.
@@ 6.2 Power logarithmic linked exposure example
An example belonging to the logarithmic linked exposure family is obtained by choosing the power function
@@
In order to achieve $b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$, which is necessary due to Proposition 4.1, the parameter $a$ has to satisfy
$a>\max_{z\in[0,1]}\left(\frac{b^{\prime}(z)^{2}}{b^{\prime\prime}(z)}-\tilde{b}(z)\right),$ (6.3)
where we set $\tilde{b}(z)=\left(1-\frac{z}{\alpha}\right)^{\delta}$. This reads in this example
@@
in order to find unimodal solutions to the MLE of the standard and extended maximization problems described in (6.1) and (6.2). The results displayed in Figure 5 are very similar to the ones obtained for the MBBEFD example. This can be confirmed by comparing Table 1 and Table 2, where most values coincide although they are actually different if we look at digits after the decimal point.
Figure 5: Power logarithmic linked exposure: densities of the random variable $Z|_{\{Z<1\}}$.
@@ 6.3 Sine logarithmic linked exposure example
Note that in this case, the function $b$ is in general neither concave, nor convex on the entire interval $[0,1]$. We claim that the condition $b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$ is fulfilled, which is necessary in order to obtain a distribution function due to Proposition 4.1. Moreover, the density for $z\in[0,1)$ reads
@@
$f^{\prime}_{Z}(z)=-\frac{\sin(\beta)+a}{\alpha\cos(\beta)}\,\frac{\alpha^{3}\cos(\alpha z+\beta)\left[-a^{2}+a\sin(\alpha z+\beta)+2\right]}{\left[\sin(\alpha z+\beta)+a\right]^{3}}.$ (6.8)
@@
Figure 6: Sine logarithmic linked exposure: densities of the random variable $Z|_{\{Z<1\}}$.
@@
in order to find unimodal solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results displayed in Figure 6 and Table 3. Although this example manages to produce rather skewed densities, see Figure 6, the fit does not seem to be accurate for this data, i.e. it is less accurate than in the first examples, and with AIC, we clearly prefer the previous models.
@@ 6.4 Quadratic exponentially linked exposure example
Let us now treat an example belonging to the exponentially linked exposure family, and given by the quadratic function
@@
This function $b$ is a smooth and strictly concave curve on $[0,1]$. We claim that the condition $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ is fulfilled for all $z\in[0,1]$, which is necessary in order to obtain a distribution function due to Proposition 5.1. Moreover, the density for $z\in[0,1)$ reads
@@
The quadratic logarithmic linked exposure example with the above parameters leads to a well-defined distribution. Moreover, the density $f_{Z}$ given in (6.9) is unimodal on $[0,1)$ if and only if $-\sqrt{-6\alpha}<\beta<-2\alpha-\sqrt{-6\alpha}$.
We use the domain
$\Theta=\Big{\{}\alpha<0,\beta\in\left(-\sqrt{-6\alpha}\,,\min\left(-2\alpha-\sqrt{-6\alpha},-\sqrt{-2\alpha}\right)\right)\Big{\}}$
in order to find unimodal solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results shown in Figure 7 and Table 4. Although a new family of exposure curves is used here, the results are close to the ones obtained with the power logarithmic linked exposure example or the MBBEFD example. In fact, in the extended problem, we obtain a slightly better model than in the MBBEFD class according to AIC. However, looking at Figure 7, this new model is still not satisfactory for this data.

### v2

@@ 3.2 Properties of the MBBEFD class
This derivative is symmetric around zero, which leads to the following result.
@@
scaled by a constant factor $(1+a)\log(1/b)>0$. It is symmetric around its mode $z^{*}$, and it decays slower than the Gaussian density. This now shows why the MBBEFD densities are not sufficient for general insurance claims modeling, because general insurance claims are typically positively skewed, which cannot be captured by the logistic density.
@@ 3.3 Extension of Bernegger’s idea using non-bell shaped densities
We extend the class of bell-shaped MBBEFD densities to a more general class of exposure curves, which allows, in particular, for skewness in their corresponding densities. We call this extended family the Bernegger class. In Section 2.1, we have started from a generic exposure curve $G:[0,1]\to[0,1]$ which is a non-decreasing, concave and twice continuously differentiable function with the normalizations $G(0)=0,\,G(1)=1$ and with $G^{\prime}(0)>0$. The MBBEFD exposure curve (3.1) can be reparametrized. Indeed, by using (3.6), we obtain for $z\in[0,1]$
@@
for parameters $b\geq 0$ and $a>-\min(1,b)$ chosen such that $G_{b,a}$ is an exposure curve; we refer to Section 3.1 of Bernegger [2]. This structure can be used to design exposure curve forms that do not have the bell-shape property of Proposition 3.3. We modify the modeling set-up (3.7) as follows. Choose a function $B:[0,1]\rightarrow{\mathbb{R}}$ that satisfies
@@
which ensures that the normalization property $G(0)=0$ and $G(1)=1$ is satisfied. The function $h$ will be denoted as the link function, whereas the function $b$ will be named the inner function, and we notice that Bernegger’s original choice was $b(z)=a+b^{z}$ and $h(x)=\log(x)$, meaning that he used a logarithmic linked exposure curve. We will first explore some examples using the same link function and then introduce the exponentially linked exposure curves, which use the link function $h(x)=\exp(x)$. We call the class of distributions induced by exposure curves of the form (3.9) the Bernegger class.
@@ 4 Logarithmic linked exposure family
The function $G$ is an exposure curve if and only if one of the following two holds:
@@
For general insurance pricing, we are interested into unimodal densities and we can thus derive the first derivative of $f_{Z}$ in order to characterize the maximum of the density
@@
Note that this derivative only exists if the third derivative of $b$ exists. The next result shows an explicit member of the Bernegger class that belongs to the logarithmic linked exposure family.
Example 4.3 (Two-parameter logistic distribution).
Consider a random variable $X$ following a two-parameter logistic distribution with density
$f_{X}(z)=\frac{e^{(z-\mu)/\sigma}}{\sigma\left(1+e^{(z-\mu)/\sigma}\right)^{2}},\quad\textrm{for}\,-\infty<z<\infty,\,-\infty<\mu<\infty,\,\sigma>0,$
and distribution
$F_{X}(z)=\frac{e^{(z-\mu)/\sigma}}{1+e^{(z-\mu)/\sigma}},\quad\textrm{for}\,-\infty<z<\infty,\,-\infty<\mu<\infty,\,\sigma>0.$
Let $d\in{\mathbb{R}}$ and $M>0$ in order to define the scaled lower-truncated and right-censored random variable
$Z=\frac{1}{M}\min\{(X-d)_{+},M\}\,|\,X>d.$ (4.5)
The distribution of $Z$ is given by
[…]
which implies
[…]
for $z\in[0,1]$. Furthermore, we have
$\int_{0}^{z}1-F_{Z}(s)\,ds=\frac{-\sigma}{M}\left(1+e^{(d-\mu)/\sigma}\right)\left[\log\left(1+e^{-(d+zM-\mu)/\sigma}\right)-\log\left(1+e^{-(d-\mu)/\sigma}\right)\right],$
for $z\in[0,1]$. This implies that the exposure curve of $Z$ is given by
$\begin{split}G(z)&=\frac{\int_{0}^{z}1-F_{Z}(s)\,ds}{\int_{0}^{1}1-F_{Z}(s)\,ds}\\ &=\frac{\log\left(1+e^{-(d+zM-\mu)/\sigma}\right)-\log\left(1+e^{-(d-\mu)/\sigma}\right)}{\log\left(1+e^{-(d+M-\mu)/\sigma}\right)-\log\left(1+e^{-(d-\mu)/\sigma}\right)},\end{split}$
for $z\in[0,1]$, which shows that the distribution of a (scaled) lower-truncated and right-censored two-parameter logistic random variable belongs to the Bernegger class with a logarithmic link function $h(x)=\log(x)$ and inner function $b(z)=1+e^{-(d+zM-\mu)/\sigma}$. $\blacksquare$
@@ 5 Exponentially linked exposure family
The function $G$ is an exposure curve if and only if one of the following two holds:
@@
As for the logarithmic linked exposure curves, one can then derive, using Theorem 2.1, a distribution function leading to an absolutely continuous density on $[0,1)$ and to a point mass in 1.
@@
Assume that a twice continuously differentiable function $b:[0,1]\to{\mathbb{R}}$ with $b(0)\neq b(1)$ fulfills condition (5.2) or (5.3). The exposure curve $G$ defined in (5.1) provides the distribution of a random variable $Z\sim F_{Z}$
@@
The first derivative of the density $f_{Z}$ given in (5.5) allows us to characterize its extrema and is given by
@@
Note that this derivative only exists if the third derivative of $b$ exists. The next example shows that lower-truncated and right-censored exponential random variables belong to the exponentially linked exposure family.
@@
Consider a total financial loss $X\sim\textrm{Exp}(\lambda)$ as well as a deductible $d>0$ and a maximal cover $M>0$. Then the scaled lower-truncated and right-censored insurance claim is given by
$Z=\frac{1}{M}\min\left\{(X-d)_{+},\,M\right\}\,|\,X>d,$
and its distribution reads as
[…]
for $z\in[0,1]$. This implies that the exposure curve of the random variable $Z$ is given by
$G(z)=\frac{\int_{0}^{z}1-F_{Z}(s)\,ds}{\int_{0}^{1}1-F_{Z}(s)\,ds}=\frac{e^{-\lambda zM}-1}{e^{-\lambda M}-1},$
for $z\in[0,1]$, which shows that the distribution of a (scaled) lower-truncated and right-censored exponential random variable belongs to the Bernegger class with an exponential link function $h(x)=\exp(x)$ and a linear inner function $b(z)=-\lambda zM$, where $M$ stands for the maximal cover. $\blacksquare$
We point out that the exponentially linked exposure family is equal to the logarithmic linked exposure family as stated in the next proposition. Thus, the choice of using one family over the other is mainly motivated by having simpler forms in the inner function $b$.
Proposition 5.4.
The logarithmic linked exposure family and the exponentially linked exposure family coincide.
6 Real dataset example
In this section, the goal is to exploit some examples belonging to the logarithmic and the exponentially linked exposure family of the Bernegger class. These examples will be used to fit general insurance claims data using the tractability of the models described in Section 2.1. In other words, MLE can directly be used since the distribution functions as well as their associated densities are of closed form.
For this, we will vary the choice of the inner function $b(z)$. In the following, $\theta\in\Theta$ will denote the set of parameters appearing in the inner function. Using a dataset for $Z$ taking values in $(0,1]$, we will fit the models with MLE, maximizing the log-likelihood function
@@ 6 Real dataset example
Min. Q1 Median Q3 Max. Mean
Normalized claims $Z_{i}$ 0.00001 0.160 0.280 0.457 1 0.339
Table 1: Summary statistics of the dataset containing 126 026126026126\,026126 026 normalized lower-truncated and right-censored claims $Z_{i}$.
Some summary statistics of the claims $Z_{i}$ are provided in Table 1 and the empirical (observed) density of the claims that are strictly smaller than $1$ is shown in Figure 3, we have an empirical mean of $0.339$ and the observed point mass in 1 is $0.034$, i.e., the insurance company pays the maximal cover on $3.4\%$ of the claims.
Figure 3: Histogram (lhs) and empirical density (rhs) of the claims $Z_{i}$, only showing the claims strictly smaller than $1$, and the point mass in $1$ is $3.4\%$.
These normalized losses $Z_{i}$, along with different models from the Bernegger class, will be used to solve the above MLE maximization problems in order to produce the results of this section. For this, we use the R function optim in order to minimize the sum of the negative of the log-likelihoods evaluated at $Z_{i}$, after having possibly transformed our parameters in a way to ensure that they lie in a suitable open domain, this is described below. The results of the best model from the Bernegger class will then be compared to lower-truncated and right-censored log-normal and gamma models at the end of this section. The first fitted model is the classical MBBEFD model of Bernegger [2].
@@ 6.1 The MBBEFD example
We have seen in Section 3.3 that the MBBEFD example belongs to the logarithmic linked exposure family and is obtained by choosing the inner function
@@
for parameters $b\geq 0$ and $a>-\min(1,b)$ chosen in a way to obtain a well-defined exposure curve. Using the parametrization in (3.6), it is possible to give the conditions under which this class of distributions leads to unimodal densities on $[0,1)$. Indeed, according to Proposition 3.2, such unimodal densities are obtained if and only if
@@
in order to find unimodal MLE solutions of the standard and extended maximization problems described in (6.1) and (6.2). In Figure 4, the empirical density of the data points strictly smaller than 1 is plotted in blue color. Since this corresponds to a true density integrating to one, we further show the conditional density of $Z|_{\{Z<1\}}$ of the fitted models, using the parameters obtained by solving the standard problem (6.1) (in green) and the extended problem (6.2) (in red). As for all the fitted models in this section, these parameters are provided in the appendix.. The point mass and the mean are shown in Table 2, as well as the log-likelihoods of the random variables $Z|_{\{Z<1\}}$ and $Z$, and the AIC scores that are computed using $\ell_{Z}$. We see that, as expected, the MLE solution of the extended problem gives better results, even if the densities obtained are not close to the empirical density. Proposition 3.3 helps us to understand why the fit is not accurate since the empirical density is skewed to the right, whereas the MBBEFD class only allows for symmetric densities described by the derivative of the logistic function.
Figure 4: The MBBEFD example: densities of the random variable $Z|_{\{Z<1\}}$.
@@
Table 2: The MBBEFD example: results.
@@ 6.2 Power logarithmic linked exposure example
Another example belonging to the logarithmic linked exposure family is obtained by choosing the power function
@@
In order to achieve $b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$, which is a necessary condition due to Proposition 4.1, the parameter $a$ has to satisfy
$a>\max_{z\in[0,1]}\left(\frac{b^{\prime}(z)^{2}}{b^{\prime\prime}(z)}-\left(1-\frac{z}{\alpha}\right)^{\delta}\right).$ (6.3)
This provides in this example
@@
in order to find unimodal solutions to the MLE of the standard and extended maximization problems described in (6.1) and (6.2). The results displayed in Figure 5 are very similar to the ones obtained for the MBBEFD example. This can be confirmed by comparing Table 2 and Table 3, where most of the values coincide, although they are actually different if we look at digits after the decimal point. We conclude that this power logarithmic linked example does not improve the fit provided by the MBBEFD example, although this model allows for skewness.
Figure 5: Power logarithmic linked exposure example: densities of the random variable $Z|_{\{Z<1\}}$.
@@ 6.3 Sine logarithmic linked exposure example
Note that in this case, the function $b$ is in general neither concave, nor convex on the entire interval $[0,1]$. We claim that $b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\geq 0$ holds for all $z\in[0,1]$, which is a necessary condition in order to obtain a distribution function due to Proposition 4.1. Moreover, the density for $z\in[0,1)$ reads as
@@
$f^{\prime}_{Z}(z)=-\frac{\sin(\beta)+a}{\cos(\beta)}\,\frac{\alpha^{2}\cos(\alpha z+\beta)\left[-a^{2}+a\sin(\alpha z+\beta)+2\right]}{\left[\sin(\alpha z+\beta)+a\right]^{3}}.$ (6.8)
@@
Figure 6: Sine logarithmic linked exposure example: densities of the random variable $Z|_{\{Z<1\}}$.
@@
in order to find unimodal solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results displayed in Figure 6 and Table 4. Although this example manages to produce rather skewed densities, see Figure 6, the fit does not seem to be accurate for this data, i.e., it is less accurate than the first examples, and we clearly prefer the previous models.
@@ 6.4 Quadratic exponentially linked exposure example
Let us now treat an example belonging to the exponentially linked exposure family, and given by the quadratic inner function
@@
This function $b$ is smooth and strictly concave on $[0,1]$. We claim that $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ holds for all $z\in[0,1]$, which is a necessary condition in order to obtain a distribution function due to Proposition 5.1. Moreover, the density for $z\in[0,1)$ is given by
@@
The quadratic exponentially linked exposure example with the above parameters leads to a well-defined distribution. Moreover, the density $f_{Z}$ given in (6.9) is unimodal on $[0,1)$ if and only if $-\sqrt{-6\alpha}<\beta<-2\alpha-\sqrt{-6\alpha}$.
Thus, we restrict to the domain
$\Theta=\Big{\{}\alpha<0,\beta\in\left(-\sqrt{-6\alpha}\,,\min\left(-2\alpha-\sqrt{-6\alpha},-\sqrt{-2\alpha}\right)\right)\Big{\}},$
in order to find unimodal solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results shown in Figure 7 and Table 5. Although a new family of exposure curves is used here, the results are close to the ones obtained with the power logarithmic linked exposure example or the MBBEFD example. In fact, in the extended problem, we obtain a slightly better model than in the MBBEFD class according to AIC. However, looking at Figure 7, this new model is again not satisfactory for this data.